A solid sphere rolls on a rough horizontal surface
with a linear speed 7 m/s collides elastically with a
fixed smooth vertical wall. Then the speed of the
sphere when it has started pure rolling in the
backward direction in m/s is.

Grade:12

2 Answers

vikas askiitian expert
13 years ago

when it collides with wall then its linear momentam get reversed but there is no change in angular speed ...

speed of center of mass is reversed ...

just after collision ,

angular momentam = mvr - Iw = -mur + 2mur/5 = -3mur/5

after a long time , total angular momentam = mvr + Iw

                                                              =mvr + 2mvr/5 = 7mvr/5

since angular mometam is conserved so

7mvr/5 = 3mur/5

v = 3u/7

u = 7 so

v = 3m/s

approve if u like my ans

Jaideep Singh Puri
11 years ago

I am not able to understand the direction of linear momentum in this question.